# Universal Vessel Segmentation for Multi-Modality Retinal Images

### v2

@@ IV-B2 Segmentation Network
An initial learning rate of 1e-3 is used to train the model for 150 epochs, then decayed to 1e-4 to train for another 250 epochs. Adam optimizer is used with a weight decay of 1e-3 and momentum of $(0.9,0.999)$. We train with a batch size of 4. The weights $\lambda_{tc}$ and $\lambda_{ts}$ are 0.05 and 0.0002, respectively. All the testing images are first padded to square and resized to 768$\times$768 to pass through the translation and the segmentation network. Then the prediction is resized back to the original image size to compute the metrics with the ground truth labels. The implementation of both image translation and the segmentation parts are under the PyTorch framework. The computation of persistent homology relies on the GUDHI library [82].
@@ IV-B3 Data Augmentation
Figure 9: Ablation studies on how each proposed step in the pipeline improves the performance of the universal vessel segmentation model.
Figure 10: Qualitative comparison for how each proposed step imporves the segmentation quality (better zoom in to view). From top to bottom: CF, MC, FA, FAF, IR modality.
@@ IV-C Comparison with Finetuned SOTA Methods
To the best of our knowledge, this work is the first universal retinal vessel segmentation model and there is no previous method that we can fairly compare with. Therefore, we compared with state-of-the-art (SOTA) methods which are finetuned individually on each of the 7 datasets in a fully-supervised fashion. Since we use three new datasets and the works in the retinal vessel segmentation literature have inconsistent dataset choice, dataset split, and inconsistent choice of evaluation metrics, we prepare a new benchmark where all the methods are evaluated in a consistent way for fair comparison. The baseline methods are run using respective recommended training details and in a five-fold experiment that splits the whole dataset by 5 groups. Each fold uses one group for testing and the other four for training. Eventually, all images are used for testing so as to make sure that the average score is computed on the same images with our universal model (which is directly tested on all the images). The baseline methods include U-Net [28], CS2-Net [30], W-Net [39], FA-Net [40], GT-DLA-dsHff [31] and WS-DMF [38]. The training and testing of the baselines are done in original image size except for the HRF dataset. The images in HRF dataset are too large to be trained on our 24GB GPUs. Therefore, we first pad the image and ground truth to square, then downsampling them to 768$\times$768 for training and testing. The predicted vessel likelihood is then upsampled to the original image size to compare with the original ground truth and compute the metrics.
@@
Since the use of evaluation metrics is inconsistent in the literature, in our work we use most of the commonly-used metrics, including the pixel-wise accuracy (Acc.), the Dice score, the specificity (Sp), the sensitivity (Se), the Precision (Pr), the F1 score and the Mathew Correlation Coefficients (MCC). In addition, since topological learning is an important part of our work and the topological accuracy is important in retinal vessel segmentation, we also use three topological metrics which are commonly used in the segmentation of more general tubular structures. These metrics include the clDice score [62] and the 0-/1-Betti errors ($\beta_{0},\beta_{1}$) [63]. All of the metrics are the larger the better except for the 0-/1- Betti errors, which are the smaller the better. For all of them reported metrics, we compute them at the original image size and averaged over all test images.
@@
The results in Table I, Fig. 8 show that our proposed model, despite allowing universal segmentation, still possesses comparable performance with SOTA finetuned methods in a general sense. However, for some modalities there still exist non-neglectable performance gap, especially for the FA. As shown by an example in the second row of Fig. 8, the FA modality (particularly from the Optos camera) has very high contrast between the vessel and the background. As a result, more ending vanishing vessels are visible on the image and the universal model often fails to identify them. The reasons are two-fold: (1) The image translation introduces non-avoidable detail loss; (2) The topological loss function makes the model more conservative in predicting a pixel as vessel.
@@
Figure 11: Comparison of the final segmentation performance between using different methods for image translation (reporting pixel-wise accuracy).
@@
Figure 12: Qualitative examples of the translated images using different image translation methods. (best zoom in to view)
IV-D Ablation Studies
IV-D1 Pipeline
@@
In this part, we study how the main components in the proposed pipeline improve the segmentation quality. All the experiments in the ablation studies adopt the same training configurations, data augmentations and network structure, unless otherwise stated. We first run a baseline experiment where no image translation nor topological learning is applied. In another word, we simply train a segmentation model on the DRIVE training set and apply the model to the multi-modality retinal images. As shown by Fig. 9 and Fig. 10, the performance of the model is very poor. After applying the image translation (CycleGAN), the segmentation quality notably improves, especially for the non-CF modalities. However, we still observe some results where the vessel boundary is not clear or adjacent vessels are mixing up (Fig. 10). After further introducing the topological learning, the model makes clean and accurate segmentation.
IV-D2 Image Translation Method
@@
As shown in Fig. 11, when comparing the final segmentation performance, CycleGAN demonstrates the best performance for most datasets. It is also observed that the CycleGAN translation results are visually more alike with the target CF images where the sharpness and contrast of the images are weaker than in the other modalities (Fig. 12). Furthermore, for FAF where the original images are very noisy, CycleGAN also performs better in reducing the noise and generating a smoother translated image.
Figure 13: Qualitative results for section IV-D3, From top to bottom: CF, MC, FA, FAF, IR.
TABLE II: Comparison between different choices of topological loss functions. Light blue-conventional pixel-wise metrics, light red-topological metrics, light green-hybrid metric.
Dataset $\mathcal{L}_{tc}$ $\mathcal{L}_{ts}$ Acc. Dice clDice $\beta_{0}$ $\beta_{1}$
$STARE$ 97.28 79.25 83.62 65.5 16.00
$\checkmark$ - 97.29 78.83 83.14 114.7 16.2
- $\checkmark$ 97.08 78.20 82.86 10.7 15.4
$\checkmark$ $\checkmark$ 97.48 80.51 85.21 35.1 12.6
$ChaseDB1$ 96.57 77.03 78.34 92.7 10.5
$\checkmark$ - 96.56 76.62 77.62 151.6 30.8
- $\checkmark$ 96.34 75.32 77.22 19.1 6.4
$\checkmark$ $\checkmark$ 96.55 77.34 78.34 70.1 7.5
$HRF$ 96.52 67.56 66.55 160.5 66.7
$\checkmark$ - 96.59 67.79 66.61 168.5 57.5
- $\checkmark$ 95.99 63.43 61.67 7.4 50.9
$\checkmark$ $\checkmark$ 96.60 68.50 68.26 70.6 55.5
$IOSTAR$ 96.48 77.51 84.61 87.6 8.0
@@
IV-D3 Topological Segmentation Loss Function
In this part we compare how the combinations of topological segmentation loss functions impact the performance of the UVSM. Table II and Fig. 13 show that, both $\mathcal{L}_{tc}$ and $\mathcal{L}_{ts}$ helps with overcoming the domain gap of the non-CF modalities. It is observed that $\mathcal{L}_{tc}$ is unable to address the inverted vessel color in FA images and tend to instead predict the boundary of the vessel, which results in a ‘sandwich’ pattern. In addition, using only $\mathcal{L}_{tc}$ adds saw artifacts and discontinuities to the segmentation. $\mathcal{L}_{ts}$ is able to address the problem of FA modality and improves the discontinuities along the vessel, however, it is observed that if only $\mathcal{L}_{ts}$ is used, the vessel boundary will be smoothed and the morphological quality of the segmentation will be poor. With a combination of $\mathcal{L}_{tc}$ and $\mathcal{L}_{ts}$, the saw and smoothing artifacts counteracts, and the challenge in predicting FA vessels is solved, producing the optimal results.
@@ V-A Limitations
As demonstrated and discussed in the experiment section, our current universal segmentation has several limitations to improve upon. Firstly, although the primary vessels are well predicted, the model tends to overprune the vessels, which leads to thin vanishing vessels appearing at the peripheral area of the retina or near the macula are missing in the segmentation. This is caused by both the non-neglectable detail loss of the image translation process and the effect of topological loss function. Secondly, our current work only addresses the vessel segmentation problem for multi-modality retinal images in a most commonly-used 30-60 degree field-of-view. Our work does not include the ultra-wide-field 135 degrees retina images. Moreover, we does not consider the commonly-used 10-15 degree OCTA image, as its imaging principle is different. These two types of images are considerably different with the conventional modalities and requires further research to be included into the universal vessel segmentation model.
@@ V-B Conclusion and Future Work
In this paper, we propose a foundational universal vessel segmentation model for multi-modality retinal images. Compared with the past literature of retinal vessel segmentation where (1) at most time only Color Fundus modality is studied and; (2) separate models need to be trained for different datasets (from different cameras) to achieve good performance, our method is able to perform robust vessel segmentation for all commonly-used modalities from all commonly-used cameras using only a single model and achieved comparable performance with SOTA finetuned methods. Although the current model still has limitations and we are working to further improve it, we believe this work is a milestone in the community of retinal image processing and understanding and marks a new direction in the research of retinal vessel segmentation. Our future work includes improving the universal vessel segmentation model to allow them to work for a wider range of field of views and to better predict the fine vanishing vessels.

### v3

@@ IV-B2 Segmentation Network
An initial learning rate of 1e-3 is used to train the model for 150 epochs, then decayed to 1e-4 to train for another 250 epochs. Adam optimizer is used with a weight decay of 1e-3 and momentum of $(0.9,0.999)$. We train with a batch size of 1. The weights $\lambda_{tc}$ and $\lambda_{ts}$ are 0.05 and 0.0002, respectively. All the testing images are first padded to square and resized to $\times768$ to pass through the translation and the segmentation network. Then the prediction is resized back to the original image size to compute the metrics with the ground truth labels. The implementation of both image translation and the segmentation parts are under the PyTorch framework. The computation of persistent homology relies on the GUDHI library [89].
@@ IV-C Comparison with Finetuned SOTA Methods
IV-C1 UVSM on retinal images in conventional field-of-view
@@
To the best of our knowledge, this work is the first universal retinal vessel segmentation model and there is no previous method that we can fairly compare with. Therefore, we compare with state-of-the-art (SOTA) methods which are finetuned individually on each dataset either by full-supervision or unsupervised domain adaptation. Since we use three new datasets and the works in the retinal vessel segmentation literature have inconsistent dataset choice, dataset split, and inconsistent choice of evaluation metrics, we prepare a new benchmark where all the methods are evaluated in a consistent way for fair comparison. The baseline methods are run using the recommended training details in respective papers and in a five-fold experiment that splits the whole dataset by 5 groups. Each fold uses one group for testing and the other four for training. Eventually, all images are used for testing so as to make sure that the average score is computed on the same images with our universal model (which is directly tested on all the images). For the unsupervised domain adaptation methods, we use the DRIVE dataset with annotation as the source domain and the per-fold training set of each dataset as the target domain for training. The supervised method baseline includes U-Net [31], CS2-Net [33], W-Net [42], FA-Net [43], GT-DLA-dsHff [34] and WS-DMF [41]. For domain adapation, the baseline includes Pse. Lab. (pseudo-label-based)[42], FRR-TSNet (teacher-student network-based) [77], MLAL (adversarial learning-based) and MLAL-PDSF (hybrid adversarial learning and teacher-student) [76]. The training and testing of the baselines are done in original image size except for the HRF dataset. The images in HRF dataset are too large to be trained on our 24GB GPUs. Therefore, we first pad the image and ground truth to square, then downsample them to $\times768$ for training and inference. The predicted vessel likelihood is then upsampled to the original image size to compare with the original ground truth and compute the metrics.
@@
Since the use of evaluation metrics is inconsistent in the literature, in our work we adopt a consistent benchmark and use most of the commonly-used metrics, including the pixel-wise accuracy (Acc.), the Dice score, the specificity (Sp), the sensitivity (Se), the Precision (Pr), the F1 score and the Mathew Correlation Coefficients (MCC). In addition, although the main goal of topological feature learning in this work is for domain generalization, topological accuracy is still important in retinal vessel segmentation. Therefore, we also use three topological metrics which are commonly used in the segmentation of more general tubular structures. These metrics include the clDice score [65] and the 0-/1-Betti errors ($\beta_{0},\beta_{1}$) [66]. For all of the reported metrics, we compute them at the original image size and average over all test images.
@@
The results in Table II, Table III and Fig. 8 show that our proposed model, not only allowing universal segmentation, but also has comparable performance with SOTA finetuned supervised methods and outperforms the finetuned unsupervised domain adaptation methods. However, for some modalities there still exist non-neglectable performance gap after the supervised methods, especially for the FA. As shown by an example in the third row of Fig. 8, the FA modality (particularly from the Optos camera) has very high contrast between the vessel and the background. As a result, more ending vanishing vessels are visible on the image and the universal model often fails to identify them. The reasons are two-fold: (1) The image translation introduces non-avoidable detail loss; (2) The topological loss function makes the model more conservative in predicting a pixel as vessel.
Figure 9: Qualitative results on Ultra-Wide-Field (UWF) retinal images. From left to right: image, translated image, UVSM segmentation, ground truth. Red arrows show the outlier patterns in the peripheral causing false segmentation, which is a current limitation of the UVSM.
Figure 10: Qualitative result on how each component in the proposed pipeline acts and further comparison with other domain adaptation methods.
Figure 11: Ablation studies on how each proposed step in the pipeline improves the performance of the universal vessel segmentation model. ‘I2I’ refers to image-to-image translation.
IV-C2 UVSM on OCTA and UWF
We separately evaluate OCTA modality and Ultra-Wide-Field retinal images (UWF) since they have considerably different field of views and the image feature in the OCTA modality is significantly different from the other modalities. In addition, vessel segmentation of these two types of retinal images has independent literature and baselines, respectively. For the metrics, we follow and use the metrics in [86] for OCTA modality and the metrics in [90] for UWF imaging.
For OCTA, we compare with SOTA supervised methods: UNet++ [91], FARGO [92], FRNet-base and FRNet [86]. We observed that the UVSM has good performance in predicting the major vessels. But the prediction for the thin ending vessels is usually missing (Fig. 8). Generally speaking, there are still notable performance gap between the UVSM and SOTA finetuned methods (Table IV). But keep in mind that we only use 19 annotated CF images to train the model, while our baseline methods are finetuned on OCTA training set with highly similar data with the testing set and are over 7 times larger. Overall, the zero-shot results on OCTA shows the strong generalization ability of our UVSM and the effectiveness of our proposed learning methods.
For UWF, we compare with SOTA supervised methods: LKM-UNet [93], Swin-Umamba [94], EM-Net [95] and Serp-Mamba [90]. Experiment shows that the UVSM has strong performance in the center area of the retina but can generate more artifacts at the boundary of the image (Fig. 9). The overall performance is comparable with the state-of-the-art methods (Table IV). The results on UWF images again shows the the strong generalization ability of our proposed model.
@@ IV-C2 UVSM on OCTA and UWF
Figure 12: Comparison of the final segmentation performance between using different methods for image translation. Reporting metric: Dice score.
@@
Figure 13: Qualitative examples of the translated images using different image translation methods. From top to bottom: MC, FA, FAF, IR, OCTA. (best zoom in to view)
IV-D Comparison with general-purpose models
Although great progress has been made in general-purpose segmentation models in recent years, they struggle with specialized image and structures such retinal image and curvilinear structures. To further highlight the need for a universal retinal vessel segmentation model, we compare with three recent foundational vision models: SAM2 [14], Grounded SAM2 [15] and the SOTA variant for general medical image segmentation-MedSAM2 [16]. For Grounded SAM2, we use a uniform prompt word ‘vessels.’. For SAM2 and MedSAM2, we provide a manual point prompt with over 100 positives points on the vessel and over 20 negative points on the background for the first image of each dataset and propagate the mask to all following images. Due to space constraint, we report the result per modality, i.e., average the score over all images (from one or multiple datasets) from each modality. The UWF image is excluded from the MC modality since the extremely high resolution of UWF images makes it infeasible to run these models on our available hardware. The result in Table V shows that our proposed UVSM has much better performance than these general-purpose segmentation models.
IV-E Ablation Studies
Due to high-volume of experimental results, for all the ablation studies, we report the result per modality the same way in section IV-D.
IV-E1 Pipeline
@@
In this part, we study how the main components in the proposed pipeline improve the segmentation quality. All the experiments in the ablation studies adopt the same training configurations, data augmentations and network structure, unless otherwise stated. We first run a baseline experiment where no image translation nor topological learning is applied. In another word, we simply train a segmentation model on the DRIVE training set and apply the model to the multi-modality retinal images. As shown by Fig. 11, the performance of the model is very poor. Both image translation and topological learning improves the performance of the model, while topological learning plays a more dominating role. Optimal performance is achieved by using both I2I and topological learning. An interesting observation is that, for the FA modality, the vessel color is inverted and the model tends to predict the boundary of the vessels as foreground rather than the vessel itself, resulting in a ‘sandwich’ pattern. None other domain adaptation method can solve this issue except for the proposed topological learning (Fig. 10).
IV-E2 Image Translation Method
@@
As shown in Fig. 12, when comparing the final segmentation performance, CycleGAN demonstrates the best performance for most datasets. It is also observed that the CycleGAN translation results are visually more alike with the target CF images where the sharpness and contrast of the images are weaker than in the other modalities (Fig. 13). Furthermore, for FAF where the original images are very noisy, CycleGAN also performs better in reducing the noise and generating a smoother translated image. Additionally, UNSB fails on the OCTA modality while StyleDiffusion and CycleGAN show better generalization ability.
Figure 14: Qualitative results for section IV-E3, From top to bottom: CF, MC, FA, FAF, IR, OCTA
TABLE VI: Comparison between different choices of topological loss functions. While the primary goal is to improve domain generalization, we still choose to report the topological metrics here since they are more relevant to the structural properties of the vessels.
Modality $\mathcal{L}_{tc}$ $\mathcal{L}_{ts}$ Acc. Dice clDice $\beta_{0}$ $\downarrow$ $\beta_{1}$ $\downarrow$
CFCF - - 96.72 73.57 74.64 114.7 35.5
$\checkmark$ - 96.75 73.43 74.31 149.8 38.5
- $\checkmark$ 96.37 71.00 72.02 12.1 27.3
$\checkmark$ $\checkmark$ 96.80 74.38 75.75 61.9 29.0
MCMC - - 96.48 77.51 84.61 87.6 8.0
@@
OCTAOCTA - - 89.50 33.09 33.51 83.2 7.9
$\checkmark$ - 94.29 62.10 64.49 102.0 6.6
- $\checkmark$ 92.22 50.89 52.45 16.8 8.3
$\checkmark$ $\checkmark$ 95.52 72.54 76.36 13.1 7.6
IV-E3 Topological Segmentation Loss Function
In this part we compare how learning different combinations of topological features impact the performance of the UVSM. Table VI and Fig. 14 show that, both $\mathcal{L}_{tc}$ and $\mathcal{L}_{ts}$ help with overcoming the domain gap of the non-CF modalities. It is observed that $\mathcal{L}_{tc}$ is unable to address the inverted vessel color in FA images and tend to instead predict the boundary of the vessel, which results in a ‘sandwich’ pattern. In addition, using only $\mathcal{L}_{tc}$ adds saw artifacts and discontinuities to the segmentation. $\mathcal{L}_{ts}$ is able to address the problem of FA modality and improves the discontinuities along the vessel, however, it is observed that if only $\mathcal{L}_{ts}$ is used, the vessel boundary will be too smoothed and the morphological quality of the segmentation will be poor. With a combination of $\mathcal{L}_{tc}$ and $\mathcal{L}_{ts}$, the saw and smoothing artifacts counteracts, and the challenge in predicting FA vessels is solved, producing the optimal results.
TABLE VII: Comparison of using different segmentation network structures for UVSM. Reporting metric: Dice score.
Backbone CF MC FA FAF IR OCTA
U-Net 74.38 78.38 75.00 75.49 74.66 72.54
FRU-Net 75.02 77.33 73.99 76.32 76.00 62.85
MCG&BSA-Net 75.53 76.94 76.23 76.34 74.18 65.99
RCARU-Net 76.02 78.00 75.70 76.07 75.28 60.27
Wave-Net 75.20 76.73 76.43 76.38 75.48 71.07
ResDoU-Net 76.01 77.89 79.13 76.15 75.49 72.71
IV-E4 Segmentation Network Structure
We also tested and compared multiple most recent network structures in retinal vessel segmentation as the segmentation model backbone for our UVSM, namely, FRUNet [77], MCG&BSA-Net [76], RCARU-Net [96], Wave-Net [97] and ResDoU-Net [98]. Table VII presents that with different backbones, universal retinal vessel segmentation is still achieved, which shows the effectiveness and flexibility of our proposed learning methods. In addition, with some more recent backbones, the overall performance of the UVSM is further improved.
@@ V-A Limitations
As demonstrated and discussed in the experiment section, our current universal segmentation has several limitations to improve upon. Firstly, although the primary vessels are well predicted, the model tends to over-prune the vessels, which leads to thin vanishing vessels appearing at the peripheral area of the retina or near the macula are missing in the segmentation. This is caused by both the non-neglectable detail loss of the image translation process and the effect of topological loss functions. The effect is prominent in the OCTA and FA modalities where finer vessels are visible. Furthermore, the robustness of the current UVSM to outlier patterns in the peripheral area of UWF images can still be improved.
@@ V-B Conclusion and Future Work
In this paper, we propose the universal retinal vessel segmentation task and the very first algorithm for this task, offering a generalizable and more practical solution to the retinal vessel segmentation problem. Compared with the past literature of retinal vessel segmentation where (1) at most time only Color Fundus modality is studied and; (2) separate models need to be finetuned for different modalities and cameras to achieve good performance, our method is able to perform modality and camera-agnostic vessel segmentation for multi-modal retinal images and achieves comparable performance with SOTA finetuned methods. Although the current model still has limitations and we are working to further improve it, we consider this work as a step forward in the community of retinal image processing. Future works include improving the universal vessel segmentation model to get better performance on fine vessels and to improve the model’s robustness to outlier patterns on the retina.